### v1

# Towards Embedding Dynamic Personas in Interactive Robots: Masquerading Animated Social Kinematics (MASK)

Abstract
This paper presents the design and development of an innovative interactive robotic system to enhance audience engagement using character-like personas. Built upon the foundations of persona-driven dialog agents, this work extends the agent’s application to the physical realm, employing robots to provide a more immersive and interactive experience. The proposed system, named the Masquerading Animated Social Kinematics (MASK), leverages an anthropomorphic robot which interacts with guests using non-verbal interactions, including facial expressions and gestures. A behavior generation system based upon a finite-state machine structure effectively conditions robotic behavior to convey distinct personas The MASK framework integrates a perception engine, a behavior selection engine, and a comprehensive action library to enable real-time, dynamic interactions with minimal human intervention in behavior design. Throughout the user subject studies, we examined whether the users could recognize the intended character in film-character-based persona conditions. We conclude by discussing the role of personas in interactive agents and the factors to consider for creating an engaging user experience.
I Introduction
Interactivity in robots [1, 2, 3] has the ability to establish meaningful connections with humans, thereby greatly improving the user experience by creating immersive relationships. Integrating character-like attributes into these robots makes interactions more human-like and relatable, enhancing emotional connections and creating memorable experiences. Interactive robots embedded with personalized personas can offer companionship for combating loneliness or deliver standout performances in the entertainment sector. Leveraging this potential, we focus on building an interactive robotic system that takes a step towards adapting flexible personas that can take on the likeness of personalities or film characters.
We hypothesize that by embedding a character template into robotic behavior, interactive agents can convincingly embody distinct personas for users to engage with. In particular, we aim to achieve this through a simple finite-state machine framework, which determines robotic actions from an observed user’s behavior, a current robot’s behavior, and a persona. We focus on non-verbal communication, where the social cues come from gestures, gazes, or facial expressions. For instance, if a robot demonstrated actions representing fear when the user is greeting it, the users would be led to associate a “shy” or “introverted” character with the robot. Furthermore, our system aims to easily adopt diverse personas, similar to a robot putting on and switching masks, requiring minimal human effort to design such robots.
To realize this vision of immersive persona-driven robots, we introduce the Masquerading Animated Social Kinetics (MASK) framework, crafted to facilitate persona-based non-verbal interactions designed to engage audiences. The system comprises the integration of three key components [1]: a perception engine that extracts meaningful features from users’ 3D body pose, a behavior selection engine that selects appropriate robotic behavior within the context, and an action library housing a collection of robot motions and facial expressions. We introduce a persona-infuser module to automate the process of generating behavior transitions with non-verbal cues and a given persona leveraging large language models (LLMs) [4], which is utilized in behavior selection engine. Notably, our framework enables minimal human intervention in behavior selection design, enabling the swift creation of interactive robotic agents imbued with distinct personas through textual input.
II Related Work
The influence of an agent’s persona on a user’s experience has been actively explored in various domains, including robotics applications [5] and dialog agents [6, 7, 8]. Several papers [5, 2, 3, 9] have focused on building personality-driven agents to provide the user’s immersive experience. Van et al. [2] focused on developing robot personalities expressed through nonverbal cues with distinct personality traits like extroversion and introversion. Mohammadi et al. [3] investigated whether a robot with a socially engaged personality is more accepted than one with a competitive personality through a case study involving a dice game. However, previous work on robotic personas was restricted to the personality with a manual design from the experts. In this work, we aim to adapt various types of personas with minimum human effort by utilizing large language models (LLMs), including personality and fictional characters, to enhance flexibility.
Prior studies [6, 7, 8] have explored the ability of language models to adapt specific personas or personalities. Safdari et al. [7] discussed the impact of synthetic personality in large language models (LLMs) on conversational agents, introducing a method for assessing and validating personality tests on LLMs and for influencing the personality in LLM-generated text. Lee et al. [8] demonstrated distinct LLM-based behaviors that enable more authentic and context-aware human-robot interactions by integrating non-verbal cues. However, the approach has limitations in the latency of requesting API calls, and the explored persona was dialog-centric rather than the character or style of the robot.
III Problem Formulation
The platform used in this work is the Ambidex [10], a tendon-driven mechanism consisting of a head, waist, and two arms. A screen which forms the head of the robot displays the robot’s facial expressions. The robot utilizes a 1 kHz low-level control system that computes the required torque for controlling the specified joint positions of the robot. A Zed2i camera [^1] https://www.stereolabs.com placed at the top of the robot detects the user’s body pose. The robot is encased within a transparent acrylic barrier to ensure interaction safety. In our setup, visitors can interact with the robot non-verbally through body language. For this work, the robot was placed in the cafe in NAVER 1784222https://www.navercorp.com/en/naver/1784, a space open to the public.
Figure 1: The proposed system architecture. The system is composed of a perception engine, a behavior selection engine, and an action library, where the action library and the behavior databases are pre-built components.
IV Proposed Method
This section describes technical details for automating the design process in a persona-based interactive system. The system’s architecture, as depicted in Figure 1, contains three main modules: a perception engine, an action library, and a behavior selection engine. At runtime, the perception engine estimates the 3D body poses of users seen in the environment via the Zed2i camera and refines human pose information to discrete observations and is used to obtain a “curiosity score”. This data is used by the behavior selection engine, which contains the state machine defining the robot’s persona-infused behavior. In the behavior selection engine, the robot selects the subsequent motion and the corresponding facial expression based on the observation and the current state. The engine’s behavior database maps observations and the robot’s persona to specific state-action transitions. This behavior database is pre-built and populated by a persona-infuser, which leverages large language models [4] to construct these transitions autonomously, alleviating the need for manual design. Finally, with the selected behavior, the robot physically displays the motion and the facial expressions that are stored within the action library.
IV-A Non-Verbal Cues
MASK employs non-verbal communication for human-robot interaction, with users interacting with the robot through body poses ($p$). As such, we examine non-verbal cues generated by non-verbal cues (presented in Table I) to drive the interaction. Human nonverbal cues (observations $o(p)$) are defined as elements that form the observation space ($\mathcal{O}$), encompassing all possible combinations of each observation element. Robots’ reactions utilize a discrete set of generated motions and facial expressions, which is the state space ($\mathcal{S}$), also presented in Table I.
Human Observation (72) # Raised Hands $\in\{0,1,2\}$,
[…]
Robot Motion (13) 1. wave hand big 2. wave hand small
3. look around 4. attract to come closer
5. small bow 6. cry 7. push away
8. hide away 9. read book 10. standstill
10. yawn 11. teasing 13. cross arms
Facial Expression (12) 1. neutral 2. smile 3. cry
4. angry 5. scared 6. excited
7. reading book 8. confused 9. yes
10. tongue out 11. yawn 12. nod head
TABLE I: Non Verbal Cues. The cues for human observation include four factors with 72 possible combinations, while robot cues contain 13 motions and 12 facial expressions with 156 possible combinations.
IV-B Automated Persona Infuser via LLMs
Figure 2: Illustration of the proposed Persona Infuser, which generates the behavior database via LLM.
In the quest to embed robots with distinct personas, we introduce a persona infuser that constructs a behavior database. This database, $\mathcal{D}$, acts as a blueprint for persona-driven behavior, encoding all possible combinations of states, observations, and transitions by large language models [4] (LLMs). The database is pre-built prior to runtime to reduce potential latency from language model inference, ensuring smooth and responsive robot actions. We utilize gpt-4-0613 model [^3] https://platform.openai.com/docs/models for our MASK  system and our experiments. The input of the LLM [4] becomes the defined state space $\mathcal{S}$, observation space $\mathcal{O}$, and the persona $x_{p}$. Conditioned with the mentioned inputs, the LLM estimates the state transitions $\mathcal{D}=\{T(s^{\prime}|s,o)\}^{s\in\mathcal{S},o\in\mathcal{O}}$. The behavior database is formulated as a dictionary. We design a hierarchical structure for generating behavior transitions as illustrated in Figure 2.
Motion Selection
We first utilize LLMs to predict the set of motions relevant to a given persona. This procedure helps the system filter out the states unrelated to the persona. The input of the LLM, denoted as $p_{\theta}$, includes the robot’s motion cues $\mathcal{M}$, the target persona $x_{p}$, and the motion selection prompt $x_{\text{motion}}$. In the motion selection prompt, we enable the LLM to select between $0.25M$ and $0.5M$ different motions, where $M=13$ is the number of motion cues. The predicted set of motions $\mathbf{m}$ is defined as $\mathbf{m}\sim p_{\theta}(\mathbf{m}|\mathcal{M},x_{p},x_{\text{motion}})$.
Expression Integration
In this step, we establish state $s$ that pairs each selected motion ($m$) with a corresponding facial expression ($f$). LLM estimates the matching facial expression conditioned on motion, iterating the estimation through the selected motions $\mathbf{m}$. The state $s$ is defined as a pair of faces and the motion, and is as follows: $s=(f,m)\quad\text{where}\quad f\sim p_{\theta}(f|m,\mathcal{F},x_{p},x_{\text{states}})$, $m\in\mathbf{m}$, $x_{\text{states}}$ is a prompt for expression integration, and $\mathcal{F}$ is a pre-defined set of facial expression cues. The set of states is denoted as $\mathbf{s}=\{(f,m)\}^{m\in\mathbf{m}}$. In addition, we utilize LLMs to define an initial state $s_{0}\sim p_{\theta}(s_{0}|\mathbf{s},x_{p},x_{\text{init}})$, where $x_{\text{init}}$ is a prompt for defining initial state. The textual inputs of the states are designed as the names of the motion to simplify the notations to LLM.
Transition Estimation
Based on the persona-based state sets $\mathbf{s}$, we estimate the state transitions $T(s^{\prime}|s,o)$ via LLMs. To infer the robot’s next state, we repeat our evaluation process for each combination of current states ($|\mathbf{s}|$) and observations ($|\mathcal{O}|$), resulting in numerous iterations. We formulate the deterministic state transition, where LLMs directly estimate the next state $s^{\prime}$ given current state $s$ and observation $o$. The input of LLM is a state set $\mathbf{s}$ from the previous paragraph, the target persona $x_{p}$, current observation $o$, current state $s$, and the state transition prompt $x_{\text{transition}}$. The transitions are defined as follows:
$T(s^{\prime}|s,o)=\begin{cases}1&\text{if}\quad s^{\prime}=s^{\prime}_{\text{LLM}}\\ 0&\text{otherwise}\end{cases}$ (1)
$s^{\prime}_{\text{LLM}}\sim p_{\theta}(s^{\prime}|s,o,\mathbf{s},x_{p},x_{\text{transition}})$ (2)
where $s\in\mathbf{s}$ and $o\in\mathcal{O}$. Again, we iterate this process to cover all observation and state spaces.
IV-C Details of MASK
Perception Engine
During the real-time deployment phase, the perception engine detects the user’s body pose to determine the user’s state. The input of the perception engine is an RGBD image, and the output is the feature vector and a curiosity score (a score that describes how ‘interesting’ a person is based on observed kinematic quantities as used in  [1]) for each person. The camera stays stationary during the interaction, and the 3D body pose of the users with tracking is estimated via ZED SDK 4.0444https://www.stereolabs.com/docs/body-tracking/. For each individual, we transform body skeleton data into five observation categories: the number of hands raised above nose level, eye gaze, distance, hand movement speed, and approaching velocity. These observations are summarized for each person as $o(p)$ in Table I. The curiosity score determines which user to interact with in multi-person scenarios, indicating how interested each user is to the robot. Following previous work [1], the curiosity score was calculated based on the observation.
Behavior Selection Engine
Decisions for state change unfolds at two key points: when changes occur in the observation or when a person of interest changes. The person of interest, which denotes a person that the robot is to interact with, is defined as a person with a maximum curiosity in the scene. Then, the next state $s^{\prime}$ is predicted by the behavior database $T(s^{\prime}|s,o(p_{in}))$, with the current state $s$ and the observation from the person of interest $o(p_{in})$. Finally, to face the user while interacting with them, we also obtain the person’s heading direction $\theta_{p_{in}}$ with respect to the robot frame. The information about the next state and the heading direction is then transferred to the action library to display the robotic action.
Action Library
The system is built upon a pre-built action library, where each state maps to specific joint trajectories and facial expressions. The action library module expresses the motion and facial expressions chosen from the behavior selection engine. Facial expressions are directly displayed on the screen as emojis. The robot’s motions are obtained from a robust motion retargeting pipeline that is described in [11] where human-we recorded the demonstration of the desired motion suitable for human-robot interaction scenarios.
V Experiments
We conducted a human-subjects study to investigate the capabilities of the proposed method in representing the persona. Our main question is whether our system enabled users to distinguish robot’s personas while considering the effects of persona-based interactions on user experience and engagement. We hypothesize that people can successfully infer the intended character with character-based persona agents. We define $67\%$ accuracy as successful in classifying the correct persona, which is four times higher than a random guess $16.7\%$.
V-A Experimental Setup
During the survey, the participants were asked to interact freely with the robot for approximately a minute. Following this interaction, participants were asked to fill out a survey (delivered in Korean). 555 The study protocol has been approved by the ethics review board of Korea University under proposal KUIRB-2024-0069-01. Four different personas, based on famous fictional movie characters, were loaded onto the robot; these included: Ebenezer Scrooge from A Christmas Carol [12], a Minion from the Despicable Me and Minions franchise [13], The Cowardly Lion from The Wizard of OZ [14], and Spock from the Star Trek franchise [15]. We recruited 108 participants (gender: M=47, F=63, ages: 20-53) to participate in this between-subjects study where the personas were randomly assigned to participants with equal numbers. Participants were asked to interact with the robot and answer three questions as in the following order: S1: List keywords that describe the robot’s personality or behavior. S2: Which of the following characters does the robot’s behavior seem to be most similar to? (multiple choice: The Cowardly Lion from The Wizard of OZ [14], A Minion from Despicable me [13], Ebenezer Scrooge from A Christmas Carol [12], Captain America from The Avengers [16], Sloth from Zootopia [17], Spock from Star Trek [15]) S3: Rate how well the robot imitates the personality of the character you selected in the previous question (Continuous Scale 00 to $100$). We have also provided a one or two-sentence summary of each character’s personality to help users understand the character mentioned in the choices for S2.
V-B Human Subject Study Results
Figure 3: Classification Confusion Matrix for the character-based persona representing the characters Cowardly Lion (LO), Minion (MM), Scrooge (SC), Spock (SS), Sloth (SZ), and Captain America (CA), respectively.
We examine our hypothesis with classification accuracy among characters based on the results of survey questions S2 and S3. The classification accuracy was reported as $76.7\%$, higher than the threshold set on mentioned hypothesis 2 ($67\%$). The confusion matrix for classification is shown in Figure 3, where we have adjusted the user’s fitting score obtained from the survey question S6. The confusion matrix $T_{ij}$ is defined as $T_{ij}=\frac{\sum_{k}\big{(}s_{f}(k)\cdot\mathds{1}(c_{k}=c_{i})\cdot\mathds{1}(g_{k}=c_{j})\big{)}}{\sum_{k}\big{(}s_{f}(k)\cdot\mathds{1}(g_{k}=c_{j})\big{)}}$, where $k$ is an index of a participants, $s_{f}(k)$ as a fitting score observed from S3, $c_{k}$ as a chosen character from S2, and $g_{k}$ as a persona shown to the participant $k$. We observed that the users could successfully identify the Cowardly Lion from The Wizard of OZ, Minions, and Scrooge characters. However, participants appeared to misclassify the character of Spock, especially between Spock ($0.44$) and Scrooge ($0.39$). Two participants who interacted with persona Spock mentioned that the robot’s firm and blunt behavior with a neutral face leads users to feel that the robot is “grumpy”, making it easily confused with Scrooge. Table II describes the top-3 mentioned keywords for each character from survey question S4. These mentioned keywords from S1 serve as an indicator that users accurately perceive the unique characteristics of persona-driven robots, which also supports our hypothesis.
Persona Keywords
Lion from Wizard of OZ Shy, Coward, Defensive
Minions Naughty, Active, Playful
Scrooge from Christmas Carol Angry, Aggressive, Dislike others
Spock from Startrek Pococurante, Blunty, Cynical
TABLE II: Top 3 mentioned keywords for each persona from survey question 3.
VI Discussions
In our user studies, we demonstrated that participants could accurately identify and differentiate the robot’s persona across character personas, involving 108 participants. Key findings include discrepancies between user expectations of warm, competent robots and their experiences, particularly with less agreeable or expressive personas, leading to challenges in interaction context comprehension. Additionally, a significant observation was the tendency of participants to anthropomorphize the robot’s actions, actively crafting narratives to understand its behavior, which suggests the importance of user interpretation in interactive robot design. However, our approach faced limitations such as a lack of diverse robotic behaviors in longer interactions, repetitive actions due to a limited number of states, and challenges in group interactions and character representation. These insights point towards the need for a broader observation space, a more complex behavior selection mechanism, and improved design for character portrayal to enhance user engagement and the overall interaction quality with interactive robots. We anticipate that this work will inspire future efforts toward expressive and convincing persona-based robotic agents that provide immersive experiences to people.